OpenStudy (anonymous):

Can someone Help Me With A Few Math Problems? Ill Medal, Fan, And write a testamonial thing xD

OpenStudy (anonymous):

i will

OpenStudy (anonymous):

whats the quistion

OpenStudy (anonymous):

whats the solution to this equation? x/2-7=15

OpenStudy (anonymous):

a.x=4

OpenStudy (anonymous):

b.x=11

OpenStudy (anonymous):

c.x=16

OpenStudy (anonymous):

d. x=44

OpenStudy (anonymous):

x in (-oo:+oo) x/2-7 = 15 // - 15 x/2-15-7 = 0 1/2*x-22 = 0 // + 22 1/2*x = 22 // : 1/2 x = 22/1/2 x = 44 x = 44

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

any other quistion's

OpenStudy (anonymous):

how did u do that so fast .-.

OpenStudy (anonymous):

im a fast typer and i just did this

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so was that it

OpenStudy (anonymous):

Point P is at the center of the circle. What is the circumference of the circle, rounded to the nearest tenth of a centimeter? A. 14.1 centimeters B. 28.3 centimeters C. 63.6 centimeters D.127.2 centimeters

OpenStudy (anonymous):

may i see the circle

OpenStudy (anonymous):

just drag the picture to a tab and copy and past the url

OpenStudy (anonymous):

or i just need the radius plz

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry idk how to put a pic xD

OpenStudy (anonymous):

ok np

OpenStudy (anonymous):

so heres the formula C=2pi*R

OpenStudy (anonymous):

so whats 2x4.5

OpenStudy (anonymous):

9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait nbo i did it wrong srry

OpenStudy (anonymous):

@ColbyMcCreery so what do i do??

OpenStudy (anonymous):

im checking brb i have to go through my notes

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

@ColbyMcCreery did u find it?

OpenStudy (anonymous):

someone plz helpp

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

srry i took so long

OpenStudy (anonymous):

are you positive?

OpenStudy (anonymous):

yes im postivtive

OpenStudy (anonymous):

i just had to eat breakfest

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

um not rlly.-.

OpenStudy (anonymous):

well was it right

OpenStudy (anonymous):

u didnt explain o.0